OpenStudy (anonymous):

A 11-oz box of cereal costs $2.64. What is the unit price?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} oz&\$\\ \hline\\ 11&2.64\\ x&1 \end{array}\implies \cfrac{11}{x}=\cfrac{2.64}{1}\)

OpenStudy (anonymous):

24 cents per ounce

OpenStudy (jdoe0001):

hmmm I got the units off =)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} oz&\$\\ \hline\\ 11&2.64\\ 1&x \end{array}\implies \cfrac{11}{1}=\cfrac{2.64}{x}\)

OpenStudy (anonymous):

Jdoe.. i don't get what the answer is

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} oz&\$\\ \hline\\ 11&2.64\\ 1&x \end{array}\implies \cfrac{11}{1}=\cfrac{2.64}{x}\implies x=\cfrac{2.64}{11}\)